OpenStudy (anonymous):

find the shortest distance between the lines x-3/1=y-5/-2=z-7/1 and x+1/7=y+1/-6=z+1/1

myininaya (myininaya):

So...I'm having trouble reading your equations.

OpenStudy (anonymous):

Equations of lines are \[\frac{ x-3 }{1 }=\frac{ y-5 }{ -2 }=\frac{ z-7 }{1 }=r1 (say) ...(1)\] and \[\frac{ x+1 }{ 7 }=\frac{ y+1 }{ -6}=\frac{ z+1 }{1 }=r2 (say) ....(2) \] Any point P on (1) is (r1+3,-2r1+5,r1+7) Any point Q on (2) is (7r2-1,-6r2-1,r2-1) Let P and Q be the shortest points, D,r's of PQ are <x2-x1,y2-y1,z2-z1>

OpenStudy (anonymous):

D.r's are <1,-2,1> D,r's of (2) are <7,-6,1 >

OpenStudy (anonymous):

use the property that shortest distance is perpendicular to both lines and find two equations in r1,r2 and solve to find r1,r2 then find P and Q \[then PQ=\sqrt{\left( x2-x1 \right)^{2}+\left( y2-y1 \right)^{2}+\left( z2-z1 \right)^{2 }}\]

myininaya (myininaya):

why is there 1 on bottom? that is weird to me.

myininaya (myininaya):

did you just do that because it was easy for you to see certain things about the problem?

OpenStudy (anonymous):

3,5,7 are the coordinates of any points on line (1) and 1,-2 and 1 are direction ratios

OpenStudy (anonymous):

@surjithayer can u plz ans. my next ques.

OpenStudy (anonymous):

eq of a line is \[\frac{ x-x1 }{ a }=\frac{ y-y1 }{ b }=\frac{ z-z1 }{ c }\] Here a,b,c are d.r's